Circular No. 9109

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2010 A2 (LINEAR) Further to IAUC 9105, the following improved orbital elements (from

MPEC 2010-A78) for this comet are like those of a minor planet in the inner part of the main belt: T = 2009 Nov. 17.248 TT, q = 1.97580 AU, e = 0.13178, ω = 125^{o}.600, Ω = 321^{o}.073, i = 5^{o}.099 (equinox 2000.0), P = 3.43 years.

J. V. Scotti, Lunar and Planetary Laboratory, reports that there is no distinct nuclear condensation on CCD images of this comet taken by R. S. McMillan with the Spacewatch 1.8-m f/2.7 reflector on Jan. 8.2 and 12.4 UT; on Jan. 8, the coma size was ∼ 8′′ × 11′′, with a sharper edge on the northern boundary, and a long, narrow tail at least 4′.6 long in p.a. 279^{o}; a faint spike extended 0′.34 in p.a. 137^{o}.

D. Jewitt, University of California at Los Angeles; and J. Annis and M. Soares-Santos, Fermilab, report observations of P/2010 A2 from Jan. 11 and 12 UT with the WYNN 3.5-m telescope on Kitt Peak. The object appears as a point-like body of red mag 23.0 ± 0.5 located 2′′.5 ± 0′′.5 east of the apex of a broad, low-surface-brightness dust tail; the latter lacks strong condensation and is more than 5′ in length (see website URL

http://www2.ess.ucla.edu/~jewitt/MBC5.html). The detected nucleus

is ≈ 150–200 m in diameter (assuming albedo 0.1), connected to the tail by an unresolved light bridge. With semimajor axis less than Jupiter’s, and Tisserand parameter 3.6, P/2010 A2 is the fifth recognized example of a main-belt comet (MBC), differing from other MBCs in being an inner-belt object (orbit consistent with membership in the Flora family). The location of the nucleus outside the tail suggests a recent impulsive origin for P/2010 A2, perhaps from a recent collision between two previously unseen minor planets, with radiation pressure driving the separation between the nucleus and the tail. In this scenario, the bridge consists of large impact-produced particles slowly separating from the nucleus under radiation pressure and Kepler shear. Future observations are encouraged to test this hypothesis.

J. Licandro, Instituto de Astrofísica de Canarias; G. P. Tozzi, Istituto Nazionale di Astrofisica, Arcetri; and T. Liimets, Nordic Optical Telescope (NOT) and Tartu Observatory, report that 5-min R- and V -band exposures obtained on Jan. 14.945–14.985 UT with the 2.5-m NOT (+ ALFOSC) shows the presence of an asteroidal object 2′′ east of (and moving at the same rate as) the uncondensed “dust swarm” of P/2010 A2, which itself is 4′ long (177000 km at the comet’s distance) and ∼ 5′′ wide in p.a. 277^{o}. Similar reports of the comet’s appearance have been received from other observers, including R. Haver and L. Buzzi.

2010 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green